OpenStudy (sarahc):

An element has three naturally occurring isotopes. Use the information below to calculate the weighted average atomic mass of the element, showing both the setup and the final answer for the calculation. Isotope Atomic Mass Percent Abundance X 1.01 u 99.984% Y 2.01 u 0.014% Z 3.02 u 0.002%

OpenStudy (sarahc):

@jim_thompson5910 hello! would u mind helping me?

jimthompson5910 (jim_thompson5910):

Are you familiar with weighted averages?

OpenStudy (sarahc):

would it be 1.01 u

jimthompson5910 (jim_thompson5910):

no

jimthompson5910 (jim_thompson5910):

The idea is to multiply each mass with it's percentage divided by 100 then you add up those products

jimthompson5910 (jim_thompson5910):

For isotope X: 1.01*99.984/100 = 1.0098384 For isotope Y: 2.01*0.014/100 = 0.0002814 I'll let you do isotope Z

OpenStudy (sarahc):

Okay so 3.02(0.002/100)

jimthompson5910 (jim_thompson5910):

yes, giving you what

OpenStudy (sarahc):

0.0000604

jimthompson5910 (jim_thompson5910):

Add that to 1.0098384 and 0.0002814 to get 1.0098384 + 0.0002814 + 0.0000604 = ???

OpenStudy (sarahc):

0.0003418

jimthompson5910 (jim_thompson5910):

1.0098384 + 0.0002814 + 0.0000604 = 1.0101802

jimthompson5910 (jim_thompson5910):

made a typo earlier before, but 1.0101802 is the answer

OpenStudy (sarahc):

Oops I knew I added that up wrong

OpenStudy (sarahc):

thank you! I understand now

jimthompson5910 (jim_thompson5910):

yw

OpenStudy (sarahc):

Is it C? The average atomic mass recorded on the periodic table for krypton is 83.80 u. This indicates that the most abundant isotope of krypton is krypton-82. krypton-83. krypton-84. krypton-85.

jimthompson5910 (jim_thompson5910):

I'm not sure. I'm not a chem expert.

jimthompson5910 (jim_thompson5910):

but if the number 83.8 refers to the numbers after the "kyrpton" then 83.8 is close to 84, so C looks good

OpenStudy (sarahc):

Okay, thats what I figured. thanks! No one on chemistry is replying to questions :(

jimthompson5910 (jim_thompson5910):

it's probably because chem is a one-time class that people forget when they pass (i know I've forgotten a lot of it and I only had to take one class) compare this to math where everyone has to take it (at least 2 years of it when I went to HS) and math is usually more universal (compared to chem)

OpenStudy (sarahc):

Oh okay. Its too bad because chemistry is not a strong subject of mine. I like math more which says something...Lol!

jimthompson5910 (jim_thompson5910):

that's great, you'll go really far with that love of math

OpenStudy (sarahc):

Do you think you could help me with this one....Which of the following sets represents a pair of isotopes? 14C and 14N Br and Br2 9B and 11B 16O and 16O2-

jimthompson5910 (jim_thompson5910):

hmm

jimthompson5910 (jim_thompson5910):

isotopes are from the same element I think

jimthompson5910 (jim_thompson5910):

so that would eliminate A

OpenStudy (sarahc):

I was thinking c

jimthompson5910 (jim_thompson5910):

maybe, B looks like it's just the same element (just in different forms), so does D

OpenStudy (sarahc):

Alrighty! :)

OpenStudy (sarahc):

The atomic and mass numbers for four different atoms are given below. Which two are isotopes of the same element? Atom A has an Atomic Number of 100 and a Mass number of 258 Atom B has an Atomic Number of 101 and a Mass number of 258 Atom C has an Atomic Number of 102 and a Mass number of 260 Atom D has an Atomic Number of 102 and a Mass number of 259 a.A and B b.B and C c.B and D d.C and D

OpenStudy (sarahc):

D?

jimthompson5910 (jim_thompson5910):

C and D sounds good, so it's choice D (lol very confusing choice of letters)